Psychology 406 Computer Assignment #1

Create you very own data set for practice:

Take 10 pennies and 10 dimes, shake them up, and throw them on the floor. Count the number of heads among the pennies and the number of heads among the dimes. Record the results on a sheet of paper. Carry out this process 15 times, recording the same results each time. Now, go into another room and repeat the same procedure. When done, your data should be recorded in the following form:

ROOM a  ROOM b
Case Pennies Dimes Case Pennies Dimes
1  x  x  16  x  x
2  x  x  17  x  x
3  x  x  18  x  x
·  ·  ·  ·  ·  ·
·  ·  ·  ·  ·  ·
·  ·  ·  ·  ·  ·
15  x  x  30  x  x

The x’s obviously refer to your own observations. Note: if your data are exactly the same as someone else’s in class, you had better start running now.

a) Create a SYSTAT file for your data that has 30 cases, and save it on your own “flash drive”. Make sure you provide a variable for “room”, which should be a character variable with the labels of “a” and “b”.

b) In your SYSTAT file mentioned in (a), provide a variable for the total number of heads for each of the 30 cases using the transformation procedures available in the editor; also, provide a variable for the difference between the number of heads for the pennies and dimes.

c) List the contents of the SYSTAT file created in (b).

d) Using the GRAPH option:

(1) Give histograms and box-plots for each of your variables.

(2) Plot the number of pennies versus the number of dimes for each room separately (use the BY command), and for the total sample. In the total sample, label the points by the room labels of “a” and “b”.

e) Using the CROSSTABS module:

(1) Provide one-way frequency tables for each variable.

(2) For the total number of heads, create a two-way frequency table of room × total number of heads.

(3) Repeat (1) and (2) for percents.

f) Using the DESCRIPTIVE STATISTICS module:

(1) Provide the basic descriptive statistics for all your variables for each room separately (use the BY command), and for the total sample.

(2) Provide stem-and-leaf diagrams for all your variables but just for the total sample.